OpenStudy (anonymous):

Which of the following is the best estimate of the area of a circle that has a radius of 4 centimeters? 12 cm2 24 cm2 48 cm2 6 cm2

OpenStudy (anonymous):

Area of circle = pi r^2 = pi (4)^2 = 16pi = ????

OpenStudy (anonymous):

math is like my worst subject i still dont understand it

OpenStudy (anonymous):

its 12?